10        20        30        40        50        60        70 t0076 1almC  STDDSPYKQAFSLFDRHGTGRIPKTSIGDLLRACGQNPTLAEITEIESTLPAE----------------- ----------------------EQ------------------QDDFKEAFLLFDRTGDAKITLSQVGDIV 80        90       100       110       120       130     140 t0076 1almC  -----------------------------VDMEQFLQVLNRPNGFDMPGDPEEFVKGFQVFDK----DAT RALGQNPTNAEINKILGNPSKEEMNAKKITFEEFLPMLQA----AANNKDQGTFEDFVEGLRVFDKEGNG 150       160       170       180       190 t0076 1almC  GMIGVGELRYVLTSLGEKLSNEEMDELLK-GVPVKDGMVNYHDFVQMILAN TVMGAELRHVLATLGEKMTEEEVEELMKGQEDSNGCINYEAFVKHIMSV--

Figure 1: Comparison of candidate 1almC-t76-hand.a2m (lines) to reference t76-1almC-GL.a2m (letters). Reference alignment has 94 matching columns, Candidate alignment has 133 matching columns. 182 were used (for template shift+target shift) to compute shift score. The shift score is 0.187217, with a min shift of 1, max shift of 39, and mean of 5.370044. Optimal subalignment has 63 matching columns, with 62 paired matches. 121 were used (for template shift+target shift) to compute shift score. Optimum shift score is 0.265385, with a min shift of 1, max shift of 2, and mean of 1.289208. Optimum achieved by removing the 71 columns indicated with dotted lines from the candidate alignment.